User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Chair: Katie

next week's chair: Patrick

Attended: Mas, Jamie, Tiege, Patrick


Absent: Jay, Lim, Lucas

General Issues

  • hoping to get Ke and Hb onto VGOS observing by the New Year
  • New people starting next year that will need training (Becca + Eloise?) maybe Simon has some students and Jamie might have people.

Handover Issues

  • R4921 Ho:
    • Antenna drives not working. Had to go into vdesk and type drvon then slew.
    • Remaining issues with rack being corrected, field system again rebooted. Hobart field system pc also full (apparently) and needed files deleting.
    • restart hobart, fix mk4 rack and receiver comms, mk5hb locked up with spectest issue (warren)
      • Most likely due to the power failure. Brett hasnt finished working on the 26m yet.
  • R1922 and R4920 Yg:
    • ifc power levels are all over the place again, this seems to be a persistent issue, I have tried all the same things I tried last time (R4920) but it does not seem to be stabilising. is this just due to increasing RFI at Yg?
      • Problem since upgrade - Jamie will contact Gino about it and update.
/home/www/auscope/opswiki/data/pages/operations/agenda2019_11_28.txt · Last modified: 2019/11/28 02:14 by Katie Vandorou